ROTATIONALLY RESOLVED SPECTRA OF THE B^{2}Π - X^{2}Π 0_{0}^{0} AND μ^{2}Σ - μ^{2}Σ 11_{1}^{1} TRANSITIONS OF C_{6}H AND C_{6}D

\underline{D. ZHAO}, M.A. HADDAD, Institute for Lasers, Life and Biophotonics Amsterdam, De Boelelaan-1081, NL 1081 HV Amsterdam, Netherlands; H. LINNARTZ, Raymond and Beverly Sackler Laboratory for Astrophysics, Leiden Observatory, Leiden University, P.O. Box 9513, NL-2300 RA Leiden, and Institute for Lasers, Life and Biophotonics Amsterdam, De Boelelaan 1081, NL-1081 HV Amsterdam, Netherlands; W. UBACHS, Institute for Lasers, Life and Biophotonics Amsterdam, De Boelelaan-1081, NL 1081 HV Amsterdam, Netherlands.

The linear carbon chain radicals have been topic of a series of spectroscopic studies. The hexatriynyl radical C_{6}H (and deuterated equivalent C_{6}D), a member of the linear C_{n}H series, has attracted renewed interest in recent years after the astronomical identification of its chemically related anion C_{6}H^{−} and its low-lying 11_{1} μ^{2}Σ vobronic state. In this talk, rotationally resolved spectra of the B^{2}Π - X^{2}Π 0_{0}^{0} and 11_{1}^{1} transitions of both C_{6}H and C_{6}D are presented. Cavity ringdown spectroscopy is used to record the spectra in direct absorption through a supersonically expanding planar plasma. The μ^{2}Σ - μ^{2}Σ 11_{1}^{1} vobronic hot bands are observed for the first time. Heavy rotational perturbations are found in the upper levels of C_{6}D. Precise spectroscopic parameters for the 11^{1} μ^{2}Σ levels of both C_{6}H and C_{6}D are determined for the first time, and the spectroscopic parameters for the B^{2}Π 0^{0} states are also improved. The Renner-Teller interaction is also discussed to estimate the excitation energies of the low-lying 11_{1} μ^{2}Σ vibronic states.